The dipole moment (μ) of HBr (a polar covalent molecule) is 0.824D (debye), and its percent ionic character is 12.2 % . Estimate the bond length of the H−Br bond in picometers.

Note that

1 D=3.34×10−30 C⋅m and

in a bond with 100% ionic character, Q=1.6×10−19 C.

Express your answer to two significant figures and include the appropriate units.
Answer #1

% ionic character = Dobs / Dionic*100

Dionic = 0.824/12.2*100 = 6.754 D

            = 6.754*3.34*10^-30 C.m

          = 2.25*10^-29 C.m

Dionic = Q*d

Q = 1.6*10^-19 C.

d = bondlength

2.25*10^-29 = 1.6*10^-19*d

d = bond length = 1.4*10^-10 m

   = 1.4*10^2 pm